AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Feb 26 23:20:13 2022
Author
*
:
Subject
*
:
<p>In this last hour we have been playing with the cryostat temperature as dictated by Nikos, in order to have the target working properly. For this, we have increased from 80K to 140K the cryostat temperature then we saw again target until we didn't. Then we went down again to 80K. After a few minutes, with a really nice target again was gone, so we increased the temperature to 120K at 00:19. At 00:27 decrease from 120K to 80K.</p> <p>Ring current 8e7 particles/s</p> <p><strong>Merger, Tape server (WS3): </strong>1580 kB/s writing ok</p> <p><strong>Leakage current (WS4): 3.27 uA </strong></p> <p><strong>DAQ (WS2): temperature checked, clock status: ok</strong><strong>, adc calibration: ok, white rabbit check: failed</strong></p> <p><strong>Online monitoring (WS 5): done</strong></p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
2022-02-27_00-26-44xytotal.png
Original size: 1920x1014
Attachment 2:
2022-02-27_00-26-22exrate.png
Original size: 1920x1014
Attachment 3:
2022-02-27_00-25-29extotal.png
Original size: 1920x1014
Attachment 4:
2022-02-27_00-25-07temp.png
Original size: 1920x1080
Attachment 5:
2022-02-27_00-24-47goodevents.png
Original size: 1920x1080
Attachment 6:
2022-02-27_00-23-42current.png
Original size: 1920x977
Attachment 7:
2022-02-27_00-31-26.png
Original size: 1920x1080
Attachment 8:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown